OpenStudy (anonymous):

( write an equation in slope-intercept form for the line that passes through 1,-1 and is parallel to y = 2x - 4

OpenStudy (anonymous):

the slope of \(=2x-4\) is \(2\) so the question can be interpreted as find the equation of the line with slope \(2\) through the point \((1,-1)\)

OpenStudy (anonymous):

use the point slope formula \[y-y_1=m(x-x_1)\] with \(m=2, x_1=1, y_1=-1\)

OpenStudy (anonymous):

i dont understand this math bs..but thanks anyways